OpenStudy (anonymous):

f(x)=(x+5)/(x+6) what is F^-1(-9)

OpenStudy (anonymous):

\(f^{-1}(-9)\)?

OpenStudy (owlcoffee):

The inverse image of F(-9)

OpenStudy (anonymous):

yx+6y=x+5 yx-x=5-6y x(y-1)=5-6y x=(5-6y)/(y-1) y=(5-6x)/(x-1)